OpenStudy (anonymous):

Dr. Sand borrowed some money to buy new furniture for her office. She paid $521.45 simple interest on a 6.5-year loan at 6.5 percent. Find the principal

hartnn (hartnn):

the formula is just I =P N R here, I = 521.45 (interest) N = 6.5 = no. of years (time period) R = 6.5% = 0.65 (rate) just plug in values and find p .

hartnn (hartnn):

can you ?

OpenStudy (anonymous):

i am working on it

OpenStudy (anonymous):

so do i mulitplie p with n and R

hartnn (hartnn):

no, I=PNR so, P = I / (NR ) = 521.45 /(0.065*6.5) =... ?

hartnn (hartnn):

and sorry, 6.5% = 0.065

OpenStudy (anonymous):

got thanks

hartnn (hartnn):

what u get P= .. ? just to verify ..

OpenStudy (anonymous):

1234.20

hartnn (hartnn):

thats correct :)

OpenStudy (anonymous):

do you have the formula for this problem In order to help pay for college, the grandparents of a child invest $4,000 in a bond that pays 10 percent interest compounded quarterly. How much money will there be in 4 years?

hartnn (hartnn):

yes, I = P (1+r/100) ^n P = 4000 r= rate = 10 n= years = 4

hartnn (hartnn):

what do you get I =... ?

OpenStudy (anonymous):

i got 5938.02

hartnn (hartnn):

4000 (1.1)^4 = 5856.4 how u got 5938 ?

OpenStudy (anonymous):

these are the answer choices give for the this question $6,051.72 B. $5,841.12 C. $6,074.15 D. $5,938.02

hartnn (hartnn):

hmm, sorry i missed the word quarterly. A = P (1+r/4*100) ^4n then you correctly get 5938.